Potassium dichromate in acidic solution is used to titrate a solution of copper(I) ions, with which it reacts according to

Cr2O72-(aq) + 6Cu+(aq)+14H3O+(aq) -- 2Cr3+(aq) + 6Cu2+(aq)+21H2O(l)


A potassium dichromate solution is prepared by dissolving 32.18 g of K2Cr2O7 in water and diluting to a total volume of 1.000 L. A total of 19.93 mL of this solution is required to reach the endpoint in a titration of a 250.0-mL sample containing Cu+(aq). Determine the concentration of Cu+ in the original solution.

________ M.
1
Expert's answer
2019-03-15T02:34:28-0400

Solution.

Cr2O72-(aq) + 6Cu+(aq)+14H3O+(aq) -- 2Cr3+(aq) + 6Cu2+(aq)+21H2O(l)

Firstly, we must find the concentration of potassium dichromate:

C = m/M*V

C (Cr2O7 2-) = 32.18/(294.19*1) = 0.109 M

n(Cr2O7 2-) = 0.109*19.93/1000 = 0.0022 mole

Secondly, we must find the normality of the potassium dichromate solution. Since, according to the reaction equation, 6 moles of monovalent copper ions interact with 1 mole of dichromate ion, we will make up the proportion:

1 mole (Cr2O7 2-) - 6 mole (Cu +)

0.00217 mole (Cr2O7 2-) - x mole (Cu +)

x = 0.01302 mole (Cu+)

C (Cu +) = n(Cu +)/V(Cu +)

C (Cu +) = 0.01302/0.25 = 0.05208 M

Answer:

C (Cu +) = 0.01302/0.25 = 0.05208 M